# Graph Neural Networks Gone Hogwild

### v1

@@ Abstract
Message passing graph neural networks (GNNs) would appear to be powerful tools to learn distributed algorithms via gradient descent, but generate catastrophically incorrect predictions when nodes update asynchronously during inference. This failure under asynchrony effectively excludes these architectures from many potential applications, such as learning local communication policies between resource-constrained agents in, e.g., robotic swarms or sensor networks. In this work we explore why this failure occurs in common GNN architectures, and identify “implicitly-defined” GNNs as a class of architectures which is provably robust to partially asynchronous “hogwild” inference, adapting convergence guarantees from work in asynchronous and distributed optimization, e.g., Bertsekas (1982); Niu et al. (2011). We then propose a novel implicitly-defined GNN architecture, which we call an energy GNN. We show that this architecture outperforms other GNNs from this class on a variety of synthetic tasks inspired by multi-agent systems, and achieves competitive performance on real-world datasets.
@@ 1 Introduction
Message-passing graph neural networks (GNNs) compute embeddings for nodes in graph-structured data by using repeated message passing within nodal neighborhoods (Hamilton, 2020; Gilmer et al., 2017; Kipf and Welling, 2017; Veličković et al., 2018). This effectively produces a deep neural network architecture whose computation graph reflects the structure of the input graph. The inherent locality in the generation of node embeddings appears at first glance to lend itself to distributed and decentralized execution, which is desirable in a variety of contexts.
Consider a group of agents (e.g., robots or “motes” in a sensor network) which need to collectively perform a task, but that might be unreliable, have limited range of communication, possess scarce local computational resources, and lack central control. GNNs are appealing as a way to learn local communication policies that solve the distributed problem, where each agent corresponds to a node in the graph and each edge to a local communication channel. One could imagine learning an algorithm that enables a swarm of robots to localize themselves, or a collection of resource-constrained edge devices to collectively estimate environmental conditions. Another application where distributed computation is attractive is in GNN inference over large graphs, where nodes or node collections are managed on distinct machines (respecting the graph connectivity). This is especially relevant for GNN deployment on resource-constrained devices. Distributed inference may also facilitate privacy in settings where nodes correspond to entities such as individuals in a social network, by enabling local inference and precluding excessive data transmission.
There is a significant obstacle in applying GNNs in these kinds of applications: distributed and decentralized computation is often asynchronous, but most GNN architectures implicitly assume synchrony in the execution of message passing at inference time. When nodes update asynchronously, messages received from neighbors may be out of date, and the effective architecture of the GNN diverges catastrophically from the training architecture; this means the output of the GNN can exhibit arbitrarily large errors. Figure 1 illustrates the issue on a toy problem. If it were possible to perform such “hogwild” inference (nodding to the asynchronous optimization work of Niu et al. (2011)), then GNNs would unlock solutions for a significantly wider set of problem domains.
@@
Figure 1: Modifications to the computation graph of an $L$-layer message-passing GNN resulting from asynchronous, distributed, per-node inference. (a) An undirected linear graph with binary features written in the node body. (b) Synchronous per-layer execution of a 3-layer GNN on the graph from (a); all nodes update at the same time, using neighbor information from the previous layer. Arrows represent network weights. For demonstration, weights are all equal to 1, and the value of a node embedding at the next layer is the sum of incoming values. (c) Asynchronous inference of the GNN from (b). Nodes update at random times and can use neighbor information corresponding to the incorrect layer, which introduces modifications to the computation graph. Gray arrows correspond to connections that were removed from the original computation graph, and red arrows are unintended connections resulting from asynchrony. (d) To demonstrate the effect of asynchrony, we show the output of the GNN varies significantly over asynchronous runs with different node update orderings.
@@
In this work, we characterize a class of “implicitly-defined” GNNs which do not exhibit this catastrophic behavior under asynchronous inference. In these GNNs, node embeddings are implicitly defined as the solution to an optimization problem (Gu et al., 2020; Liu et al., 2021; Yang et al., 2021; Zhu et al., 2021; Ma et al., 2021; Scarselli et al., 2009). This is in contrast to GNNs in which node embeddings are computed as the output of a specific forward computation, such as graph attention networks (GAT, Veličković et al. (2018)) and graph convolutional networks (GCN, Kipf and Welling (2017)); we refer to these architectures as “explicitly-defined”. Implicitly-defined GNNs are provably robust to asynchrony given lenient assumptions on staleness and per-node update frequency, making them ideally suited to distributed, asynchronous inference. However, the diversity of existing implicitly-defined GNN architectures is limited. We contribute a novel architecture, which we call energy GNN, which exposes a rich parameterization of optimization problems by using input-convex neural networks defined on nodal neighborhoods. We show that energy GNNs outperform other implicitly-defined GNNs on a variety of synthetic tasks motivated by problems which are relevant to multi-agent systems, where synchronous inference may be undesirable. We also achieve competitive performance on tasks with benchmark graph datasets, certifying the merit of our approach even as a stand-alone GNN architecture.
Our contributions are as follows:
  • We present the first unified framework for GNNs which can execute inference asynchronously. We refer to this class of GNNs as being implicitly defined, and articulate sufficient conditions that such architectures must satisfy. We then contextualize all major GNN architectures and identify which are implicitly defined, and which are not.
  • We propose a novel implicitly defined architecture called the energy GNN. Unlike existing implicitly-defined GNNs, energy GNNs 1) incorporate edge features, a crucial property in many applications, 2) enable the generation of neighbor-specific messages, 3) enable the utilization of neighborhood attention, and 4) use neural networks for generation of messages and embeddings.
  • We provide conditions under which there are guarantees for the convergence of our architecture under partially asynchronous inference, adapting classical results from the distributed optimization literature.
  • We present several empirical studies which span synthetic datasets crafted to tease out important distinctions between GNNs which are/aren’t amenable to asynchronous inference, and several widely used GNN benchmarks to assess performance in a cross-referenced manner with existing results published in the literature.
@@ 2.1 Partially Asychronous Algorithms
Computational models for asynchronous algorithms vary depending on the constraints imposed on the sequencing or frequency of computation or communication. We consider partial asynchronism as defined by Bertsekas and Tsitsiklis (1989), which, informally, places bounds on two key characteristics: the time between updates for each node, and the amount by which data retained at any node can be out of date from the perspective of some other node(s).
Concretely, consider a collection of $n$ nodes, which could be a set of processors carrying out a distributed computation. Each node has hidden state denoted ${{\bm{h}}_{i}\in\mathbb{R}^{k}}$, which corresponds to one “block” (row) of the aggregate state ${{\bm{H}}:=({\bm{h}}_{1},\ldots,{\bm{h}}_{n})^{T}\in\mathbb{R}^{n\times k}}$. The distributed computation we aim to execute is an asynchronous iterative algorithm, where each node $i=1,2,\dots,n$ iteratively updates its state according to ${\bm{h}}_{i}:=f^{i}({\bm{H}})$ using some node-specific update function $f^{i}:\mathbb{R}^{n\times k}\to\mathbb{R}^{k}$.
Asynchrony in the execution of the algorithm is accounted for as follows. We are given a set $T^{i}\subseteq\{0,1,2,\dots\}$ of times at which each node $i$ is updated. Additionally, for each ${t\in T^{i}}$ we are given variables $0\leq\tau_{j}^{i}(t)\leq t$ which represent the time of node $j$’s last update from the perspective of node $i$, at time $t$. Note that $\tau^{i}_{j}(t)\in T^{j}$. The quantities ${t-\tau_{j}^{i}(t)\in[0,t]}$ can then be interpreted as the amount (in time) by which information associated with node $j$ is outdated or “stale” when used in the update of ${\bm{h}}_{i}$ at time $t$. For instance, the state ${T^{3}=\{0,4,7\}}$, ${T^{1}=\{1,2,3\}}$, ${\tau_{1}^{3}(4)=2}$ means that when node 3 executes its second update (at ${t=4}$), it uses the stale value of node 1 that was computed at ${t=2}$, rather than the most recent value at ${t=3}$.
@@ 2.2 Graph Neural Networks
In this section, we provide a brief overview of GNNs. Consider a directed graph with a collection of $n$ vertices ${V=\{1,...,n\}}$, and edges ${E\subseteq V\times V}$. The connectivity of the graph is contained in its adjacency matrix ${\bm{A}}\in\{0,1\}^{n\times n}$, where ${\bm{A}}_{i,j}=1$ if there is an edge from node $i$ to node $j$, and 00 otherwise. The graph may also have associated node and edge features ${{\bm{X}}\in\mathbb{R}^{n\times p}}$ and ${{\bm{E}}\in\mathbb{R}^{|E|\times r}}$, so we use ${\mathcal{G}}=({\bm{A}},{\bm{X}},{\bm{E}})$ to denote the graph (and its features). We use $\mathcal{N}(i)$ to denote the set of neighbors of node $i$, and $|{\mathcal{G}}|$ to denote the number of nodes in a graph ${\mathcal{G}}$.
We are given a dataset ${\mathcal{D}=\{({\mathcal{G}},{\bm{Y}})^{q}\}_{q=1}^{|\mathcal{D}|}}$, where each ${\mathcal{G}}^{q}$ is associated with a node-level prediction target ${{\bm{Y}}^{q}\in\mathbb{R}^{n_{q}\times J}}$, with $n_{q}=|{\mathcal{G}}^{q}|$. We focus on node-level predictions, since they often underlie graph-level predictions.
In a GNN, each node $i=1,\dots,n$ is associated with a $k$-vector embedding ${\bm{h}}_{i}$ which is updated through iterations (or layers) of message passing; we use ${\bm{H}}:=({\bm{h}}_{1},\dots,{\bm{h}}_{n})^{T}\in\mathbb{R}^{n\times k}$ to denote the aggregate embedding of nodes in the graph. The dimensionality $k$ can vary across layers but, without loss of generality, we assume in this work that it is fixed. Each layer ${\ell\in\{1,...,L\}}$ in a GNN defines a parameterized embedding function $f_{\theta}^{\ell}:{\mathcal{G}}_{i}\times\mathbb{R}^{|{\mathcal{G}}_{i}|\times k}\rightarrow\mathbb{R}^{k}$. $f_{\theta}^{\ell}$ is applied independently for each node $i=1,\dots,n$ and takes as argument a subgraph ${\mathcal{G}}_{i}\subseteq{\mathcal{G}}$ corresponding to the neighborhood of node $i$ and the subgraph’s node embeddings from the previous layer ${\bm{H}}^{\ell-1}\in\mathbb{R}^{|{\mathcal{G}}_{i}|\times k}$, and returns an embedding ${\bm{h}}^{\ell}_{i}$ for node $i$. The new embeddings ${\bm{H}}^{\ell}$ are used in the next layer. Typically, ${\bm{H}}^{0}$ is initialized as the node features ${\bm{X}}$. For convenience, we omit the layer index $\ell$ from $f_{\theta}^{\ell}$ in contexts where there is only one parameterized layer, or when referring generally to a message passing layer.
A layer can be described as performing message passing, where each node $i$ constructs a “message” ${\bm{m}}_{i}$ by aggregating information (e.g., embeddings from the previous layer and node/edge features) from nodes $j\in\mathcal{N}(i)$ in its local neighborhood, then uses this aggregate message to update its embedding. In the final layer, a readout function ${o_{\phi}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{J}}$ is applied to each embedding, which results in node predictions ${\hat{{\bm{Y}}}=(o_{\phi}({\bm{h}}^{L}_{1}),...,o_{\phi}({\bm{h}}^{L}_{n}))^{T}\in\mathbb{R}^{n\times J}}$.
Different GNN architectures are distinguished by the parameterization of the update function $f_{\theta}$. In Appendix A, we describe GAT and GCN, two widely-used architectures.
@@
This failure motivates us to distinguish between two types of GNN architectures: explicitly-defined GNNs which specify a specific feed-forward layer-wise computation, and implicitly-defined GNNs in which the layer-wise message passing updates correspond to iterations toward a fixed point. We further subdivide implicitly-defined GNNs into two types: fixed-point GNNs and the important special case of optimization-based GNNs. Explicitly-defined GNNs are widely-used but are susceptible to failure under asynchrony, while implicitly-defined GNNs are robust (as we discuss in Section 4).
@@ 3.1 Fixed-point GNNs
Convergence to a unique fixed point is guaranteed provided the update function $F_{\theta}$ is a contraction map with respect to the embeddings, i.e. $||F_{\theta}({\bm{h}})-F_{\theta}({\bm{h}}^{\prime})||\leq\mu||{\bm{h}}-{\bm{h}}^{\prime}||$ holds for $10<\mu<$ 1 and $\forall{\bm{h}},{\bm{h}}^{\prime}\in\mathbb{R}^{nk}$. Since it is difficult to design non-trivial parameterizations of $F_{\theta}$ which are contractive by construction, existing fixed-point GNN architectures (e.g. IGNN (Gu et al., 2020), EIGNN (Liu et al., 2021), and APPNP (Gasteiger et al., 2019)) are limited in diversity. These architectures use linear transformations of the node embeddings to form messages ${\bm{m}}_{i}$, where the parameters can easily be constrained such that $F_{\theta}$ is contractive. Gu et al. (2020) apply a component-wise non-expansive nonlinearity to compute updated node embeddings from these messages, while Liu et al. (2021); Gasteiger et al. (2019) take ${\bm{m}}_{i}$ as the node embeddings directly.

### v2

@@ Abstract
Graph neural networks (GNNs) appear to be powerful tools to learn state representations for agents in distributed, decentralized multi-agent systems, but generate catastrophically incorrect predictions when nodes update asynchronously during inference. This failure under asynchrony effectively excludes these architectures from many potential applications where synchrony is difficult or impossible to enforce, e.g., robotic swarms or sensor networks. In this work we identify “implicitly-defined” GNNs as a class of architectures which is provably robust to asynchronous “hogwild” inference, adapting convergence guarantees from work in asynchronous and distributed optimization. We then propose a novel implicitly-defined GNN architecture, which we call an energy GNN. We show that this architecture outperforms other GNNs from this class on a variety of synthetic tasks inspired by multi-agent systems.
@@ 1 Introduction
Coordination and control of distributed and decentralized multi-agent systems is a major project spanning engineering and computational science. Success in this program has implications across a broad spectrum of applications; autonomous vehicle navigation, energy/resource management and distribution in smart grids, and robotic swarms (for exploration, search and rescue, environmental monitoring, and construction), to name a few. Agents in these systems must take action based on direct observation and communication with the collective. At the level of any individual, a controller takes as input the agent’s ‘state representation’, which ideally unifies direct measurements and peer-derived messages in a coherent manner.
In the distributed/decentralized regime, the communication constraints associated with the system can be encoded as a graph in which each node is associated with an agent, and each edge with an agent-to-agent communication link. Adopting the graph view of the system, graph neural networks (in particular, message-passing variants (Gilmer et al., 2017; Hamilton, 2020)) have been widely studied as methods for deriving flexible, data-dependent, and parametric state representations. GNN-based state representations have been explored in a variety of applications; for example, flock formation, target tracking, path planning, goal assignment, and channel allocation in wireless networks (Jiang and Lu, 2018; Nakashima et al., 2019; Li et al., 2019; Khan et al., 2019; Jiang et al., 2020; Gama et al., 2020; Blumenkamp et al., 2021; Grattarola et al., 2021; Zhou et al., 2021; Gosrich et al., 2022; Jiang et al., 2023; Goarin and Loianno, 2024).
The motivation to use GNNs can be traced to three core features. First, a GNN is a parametric family of functions for deriving a state representation, enabling domain-specificity by choosing from this family and optimizing the state representation for a given task end-to-end. Second, message passing within a given GNN layer satisfies the constraint that agents must derive state representations using only information obtained through local communication with neighbors. Third, GNNs with multiple layers allows state representations to depend on information outside of a given agent’s local neighborhood.
All that said, there is a major conundrum in applying GNNs to these problems. On the one hand, asynchronous execution and unreliable communication are ubiquitous in real-world decentralized/distributed multi-agent systems. On the other hand, these features render conventional GNNs inoperable in the multi-layer regime, the very setting in which GNNs offer a putative advantage through longer range communication/coordination. This is because GNN architectures implicitly assume that a synchronization barrier is enforced across all nodes between layers of message passing. If nodes update asynchronously or if messages can be delayed or lost, embeddings from neighbors do not necessarily correspond with the intended layer in the GNN (from the perspective of the receiving node). The effective architecture (and therefore output) of the GNN diverges catastrophically from that used in training. Figure 1 illustrates the issue on a toy problem. Unfortunately, enforcing synchronization in a distributed and decentralized system is costly (e.g. throughput is limited by the slowest node) and is particularly difficult to enforce in systems where nodes can join or leave the network, or are prone to failure. Existing work on using GNNs for computing agent state embeddings either use single layer GNNs (where synchronization is unnecessary), or ignore the constraint of asynchronous execution altogether, limiting the prospect of fielding the method. If it were possible to perform “hogwild” inference in multi-layer GNNs (nodding to the asynchronous optimization work of Niu et al. (2011)), GNNs would be a significantly more powerful tool in computing state embeddings for decentralized multi-agent systems.
@@
Figure 1: Modifications to the computation graph of an $L$-layer message-passing GNN resulting from asynchronous, distributed, per-node inference with communication delays. (a) An undirected linear graph with binary features written in the node body. (b) Synchronous per-layer execution of a 3-layer GNN on the graph from (a); all nodes update at the same time, using neighbor information from the previous layer. Arrows represent network weights. For demonstration, weights are all equal to 1, and the value of a node embedding at the next layer is the sum of incoming values. (c) Asynchronous inference of the GNN from (b). Nodes update at random times and can use neighbor information corresponding to the incorrect layer, which introduces modifications to the computation graph. Gray arrows correspond to connections that were removed from the original computation graph, and red arrows are unintended connections resulting from asynchrony and/or message delays. (d) To demonstrate the effect of asynchrony, we show the output of the GNN varies significantly over asynchronous runs with different node update orderings.
@@
In this work we present a unified framework for GNNs which can execute inference asynchronously, adapting classical results from the distributed optimization literature. Our analysis illuminates the boundary between GNNs which are/are not amenable to asynchronous inference, and we articulate sufficient conditions for a given architecture’s robustness to asynchrony. We call the class of GNNs which are provably robust to asynchronous inference ‘implicitly-defined’; in an implicitly-defined GNN, node embeddings correspond with a solution to an optimization problem (Gu et al., 2020; Liu et al., 2021; Yang et al., 2021; Zhu et al., 2021; Ma et al., 2021; Scarselli et al., 2009). This is in contrast to GNNs in which node embeddings are computed as the output of a statically defined feed-forward computation, such as graph attention networks (GAT, Veličković et al. (2018)) and graph convolutional networks (GCN, Kipf and Welling (2017)); we refer to these architectures as ‘explicitly-defined’. In addition to an analysis of GNNs under asynchronous execution, we contribute a novel asynchronous-capable (implicitly-defined) architecture, which we call energy GNN. Our architecture exposes a rich parameterization of optimization problems using input-convex neural networks operating over nodal neighborhoods. We show that energy GNNs outperform other implicitly-defined GNNs on a variety of synthetic tasks motivated by problems which are relevant to multi-agent systems, where synchronous inference may be undesirable. We also achieve competitive performance on tasks with benchmark graph datasets, certifying the merit of our approach even as a stand-alone GNN architecture.
@@ 2.1 Partially Asychronous Algorithms
Computational models for asynchronous algorithms vary depending on the constraints imposed on the sequencing or frequency of computation or communication. We consider partial asynchronism as defined by Bertsekas and Tsitsiklis (1989), which we summarize here.
Consider a collection of $n$ nodes carrying out a distributed computation. Each node has hidden state denoted ${{\bm{h}}_{i}\in\mathbb{R}^{k}}$, which corresponds to one “block” (row) of the aggregate state ${{\bm{H}}:=({\bm{h}}_{1},\ldots,{\bm{h}}_{n})^{T}\in\mathbb{R}^{n\times k}}$. The algorithm being executed consists of iterative node updates, where each node $i=1,2,\dots,n$ iteratively updates its state according to ${\bm{h}}_{i}:=f^{i}({\bm{H}})$ using some node-specific update function $f^{i}:\mathbb{R}^{n\times k}\to\mathbb{R}^{k}$.
We are given a set $T^{i}\subseteq\{0,1,2,\dots\}$ of times at which each node $i$ is updated; this accounts for asynchrony in the execution of the algorithm. Additionally, for each ${t\in T^{i}}$ we are given variables $0\leq\tau_{j}^{i}(t)\leq t$ which represent the time associated with node $i$’s view of node $j$ at time $t$. If communication is instantaneous and reliable, then $\tau^{i}_{j}(t)$ is always equal to $t$, as node $i$’s view of all other nodes is the actual value those nodes hold at time $t$. Staleness (i.e. $\tau_{j}^{i}(t)<t$) results from delays or losses in communication between nodes. For instance, suppose ${T^{3}=\{0,4,7\}}$, ${T^{1}=\{2,3,5\}}$, and ${\tau_{1}^{3}(4)=2}$. This means that when node 3 executes its second update (at ${t=4}$), it uses the stale value of node 1 that was computed at ${t=2}$ rather than its actual value at $t=4$ (due to loss or delay of the message containing the value of node 1 which was computed at ${t=3}$).
@@
Informally, (a) states that each node updates at least once every $B$ time units, (b) states that information from other nodes can be stale by at most $B$ time units, and (c) states that node $i$ maintains the current value of ${\bm{h}}_{i}$.
@@ 2.2 Graph Neural Networks
Consider a directed graph with a collection of $n$ vertices ${V=\{1,...,n\}}$, and edges ${E\subseteq V\times V}$. The connectivity of the graph is contained in its adjacency matrix ${\bm{A}}\in\{0,1\}^{n\times n}$, where ${\bm{A}}_{i,j}=1$ if there is an edge from node $i$ to node $j$, and 00 otherwise. The graph may also have associated node and edge features ${{\bm{X}}\in\mathbb{R}^{n\times p}}$ and ${{\bm{E}}\in\mathbb{R}^{|E|\times r}}$, so we use ${\mathcal{G}}=({\bm{A}},{\bm{X}},{\bm{E}})$ to denote the graph. We use $\mathcal{N}(i)$ to denote the set of neighbors of node $i$.
In a GNN, each node $i$ is associated with a $k$-vector embedding ${\bm{h}}_{i}$ which is updated through iterations (or layers) of message passing. Each node constructs a “message” ${\bm{m}}_{i}$ by aggregating information from nodes $j\in\mathcal{N}(i)$ in its local neighborhood, then uses this aggregate message to update its embedding.
The node update equation for all nodes $i$ and layers ${\ell\in\{0,...,L-1\}}$ is as follows:
${\bm{h}}_{i}^{\ell+1}$ $=f_{\theta}^{\ell}\left({\bm{h}}_{i}^{\ell},{\bm{x}}_{i},\{{\bm{h}}_{j}^{\ell},{\bm{x}}_{j},{\bm{e}}_{ij}|j\in\mathcal{N}(i)\}\right)=u^{\ell}_{\theta}\left({\bm{h}}_{i}^{\ell},{\bm{x}}_{i},\bigoplus_{j\in\mathcal{N}(i)}m^{\ell}_{\theta}({\bm{h}}_{i}^{\ell},{\bm{h}}_{j}^{\ell},{\bm{x}}_{i},{\bm{x}}_{j},{\bm{e}}_{ij})\right)=u^{\ell}_{\theta}\left({\bm{h}}_{i}^{\ell},{\bm{x}}_{i},{\bm{m}}_{i}^{\ell}\right),$ (3)
where $\bigoplus$ is an aggregation function, and $u_{\theta}$ and $m_{\theta}$ are, e.g., neural networks. We use ${\bm{H}}:=({\bm{h}}_{1},\dots,{\bm{h}}_{n})^{T}\in\mathbb{R}^{n\times k}$ to denote the aggregate embedding of nodes in the graph. ${\bm{H}}^{0}$ is typically initialized as the node features ${\bm{X}}$.
For convenience, we omit the layer index $\ell$ from $f_{\theta}^{\ell}$ in contexts where there is only one parameterized layer, or when referring generally to a message passing layer.
In the final layer, a readout function ${o_{\phi}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{J}}$ is applied to each embedding, which results in node predictions ${\hat{{\bm{Y}}}=(o_{\phi}({\bm{h}}^{L}_{1}),...,o_{\phi}({\bm{h}}^{L}_{n}))^{T}\in\mathbb{R}^{n\times J}}$.
@@
This failure motivates us to distinguish between two types of GNN architectures: explicitly-defined GNNs which specify a specific feed-forward layer-wise computation, and implicitly-defined GNNs in which the layer-wise message passing updates correspond to iterations toward a fixed point. We further subdivide implicitly-defined GNNs into two types: fixed-point GNNs and the important special case of optimization-based GNNs. Explicitly-defined GNNs are susceptible to failure under asynchrony, while implicitly-defined GNNs are robust (as we discuss in Section 4).
@@ 3.1 Fixed-point GNNs
Convergence to a unique fixed point is guaranteed provided the update function $F_{\theta}$ is a contraction map with respect to the embeddings, i.e. $||F_{\theta}({\bm{h}})-F_{\theta}({\bm{h}}^{\prime})||\leq\mu||{\bm{h}}-{\bm{h}}^{\prime}||$ holds for $10<\mu<$ 1 and $\forall{\bm{h}},{\bm{h}}^{\prime}\in\mathbb{R}^{nk}$. Since it is difficult to design non-trivial parameterizations of $F_{\theta}$ which are contractive by construction, existing fixed-point GNN architectures (e.g. IGNN (Gu et al., 2020), EIGNN (Liu et al., 2021), and APPNP (Gasteiger et al., 2019)) are limited in diversity. These architectures use linear transformations of the node embeddings in computing messages ${\bm{m}}_{i}$, where the parameters can easily be constrained such that $F_{\theta}$ is contractive. Gu et al. (2020) apply a component-wise non-expansive nonlinearity to compute updated node embeddings from these messages, while Liu et al. (2021); Gasteiger et al. (2019) directly take ${\bm{m}}_{i}$ as the node embeddings for the next iteration.